# Limit theorems for a strongly irreducible product of independent random matrices under optimal moment assumptions

### v2

@@ 2.2. Local-to-global properties for the alignment
Let $f,g,h$ be non-zero matrices such that the product $fgh$ is well defined and let $0<\varepsilon\leq 1$. Assume that $\sigma(g)\leq\varepsilon^{2}/4$ and that $f\mathbb{A}^{\varepsilon}g\mathbb{A}^{\frac{\varepsilon}{2}}h$. Then $f\mathbb{A}^{\frac{\varepsilon}{2}}gh$.
@@
Let $f$, $g$ and $h$ be non-zero matrices such that the product $fgh$ is well defined and let $0<\varepsilon\leq 1$. Assume that $f\mathbb{A}^{\varepsilon}g\mathbb{A}^{\varepsilon}h$ and that $\sigma(g)\leq\varepsilon^{4}/4$. Then $\sigma(fgh)\leq 4\varepsilon^{-4}\sigma(f)\sigma(g)\sigma(h)\geq 4\varepsilon^{-4}\sigma(g)$.
@@
Let $12\leq k\leq n-$ 1. First by Lemma 2.9 applied to $(g_{k},\dots,g_{n})$, we have $g_{k}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k+1}\cdots g_{n}$. Again by Lemma 2.9 applied to $(^{t}g_{k-1},\dots,^{t}g_{0})$, we have $g_{0}\cdots g_{k-2}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k-1}$. Moreover, we have $g_{k-1}\mathbb{A}^{\varepsilon}g_{k}$ an $\sigma(g_{k-1}),\sigma(g_{k})\leq\varepsilon^{2}/12$. Therefore, by Lemma 2.14, we have $g_{0}\cdots g_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k}\cdots g_{n}$. ∎
@@ 2.3. Link between singular values and eigenvalues
To get (34), we apply Corollary 2.13. We get a line $l^{\infty}$ such that for any sequence $u_{n}\in U^{1}(g^{n})\setminus\{0\}$, we have $[u_{n}]\to l^{\infty}$. Let $l^{\infty}$ be such a line and $(u_{n})_{n\geq 0}$ be such a sequence. Moreover, for all $u\in U^{1}(g)$, we have $\mathrm{d}([u],l^{\infty})\leq\frac{2\sigma(g)}{\varepsilon}$. Now we only need to show that $l^{\infty}=E^{+}(g)$. Let $e\in E^{+}(g)$. Then we have $\|ge\|=\rho_{1}(g)\|e\|$ so by (32), $e\in V^{\frac{\varepsilon}{2}}(g)$. Moreover, $e$ is an eigenvector so $e\in\mathbb{K}ge$ and as a consequence $e\in U^{\frac{\varepsilon}{2}}(g)$. This reasoning holds for all powers of $g$ so we have $e\in U^{\frac{\varepsilon}{2}}(g^{n})$ for all $n$. Moreover, $\sigma(g^{n})\to 0$ by(33) so by Lemma 2.5, the projective diameter of $U^{\frac{\varepsilon}{2}}(g^{n})$ goes to zero so $[u_{n}]\to[e]$ and therefore $l^{\infty}=[e]$. ∎
@@ 3. Random products and extractions
We use the following notations. Given $\Gamma$ a metric semi-group with a unit $\mathbf{1}_{\Gamma}$, we write $\widetilde{\Gamma}:=\bigsqcup_{n\geq 0}\Gamma^{n}$ for the set of words on the alphabet $\Gamma$ and write $\odot$ for the concatenation product. We write ()()( ) for the empty word.
@@
Given $(w_{k})_{k\geq 0}\in\mathbb{N}^{\mathbb{N}}$, we write $\overline{w}_{k}:=w_{0}+\cdots+w_{k-1}$ for all $k\in\mathbb{N}$ and given $(\gamma_{n})_{n\geq 0}\in\Gamma^{\mathbb{N}}$, we write $\gamma^{w}_{k}:=\gamma_{\overline{w}_{k}}\cdots\gamma_{\overline{w}_{k+1}-1}$ and $\widetilde{\gamma}^{w}_{k}:=\left(\gamma_{\overline{w}_{k}},\dots,\gamma_{\overline{w}_{k+1}-1}\right)$ for all $k\in\mathbb{N}$. That way, for all sequence $(w_{k})_{k\geq 0}\in\mathbb{N}^{N}N$ that is not stationary to 00 and for all sequence $(\gamma_{n})_{n\geq 0}\in\Gamma^{\mathbb{N}}$, we have $\bigodot_{k=0}^{+\infty}\widetilde{\gamma}^{w}_{k}=(\gamma_{n})_{n\geq 0}$.
@@ 3.2. Rank one boundary of a semi-group
$\partial_{u}\Gamma$ :=$:=\left\{[hv]\,\middle|\,[h]\in\partial\Gamma,\;v\in E\setminus\ker(h)\right\}\subset\mathrm{P}(E)$
$\partial_{w}\Gamma$ :=$:=\left\{[fh]\,\middle|\,[h]\in\partial\Gamma,\;f\in E^{*}\setminus\ker({{}^{t}h})\right\}\subset\mathrm{P}(E^{*}).$
@@
Note that the space $\mathrm{P}(\mathrm{End}(E))$ is metrizable so the closure is characterized by sequences. Let $\pi$ be a rank one endomorphism. Saying that $[\pi]\in\partial\Gamma$ is equivalent to saying that there is a sequence $(\gamma_{n})\in\left(\Gamma\setminus\{0\}\right)^{\mathbb{N}}$ such that $[\gamma_{n}]\to[\pi]$. Note also that the product map is continuous so $\overline{\Gamma}$ is a semi-group. Therefore, for all $[\pi_{1}],[\pi_{2}]\in\partial\Gamma$, and for all $g\in\Gamma$ such that $\pi_{1}g\pi_{2}\neq 0$, we have $[\pi_{1}g\pi_{2}]\in\partial\Gamma$.
Let $v_{1}\in\partial_{u}\Gamma$ and let $f_{2}\in\partial_{v}\Gamma$. Let $v_{2}\in E$ and $f_{1}\in E^{*}$ be such that $[v_{1}f_{1}]\in\partial\Gamma$ and $[v_{2}f_{2}]\in\partial\Gamma$. By definition of the irreducibility, there is an element $g\in\Gamma$ such that $f_{1}gv_{2}\neq 0$. Let $g$ be such an element. Then we have $v_{1}f_{1}gv_{2}f_{2}\neq 0$ hence $[v_{1}f_{1}gv_{2}f_{2}]\in\partial\Gamma$. Moreover $[v_{1}f_{1}gv_{2}f_{2}]=[v_{1}f_{2}]$ because $f_{1}gv_{2}$ is a scalar, therefore $[v_{1}f_{2}]\in\partial\Gamma$. ∎
@@
Reciprocally, let $[\pi]\in\partial\nu$. By irreducibility, there exists $g\in\overline{\Gamma}_{\nu}$ such that $g(\mathrm{im}(\pi))\not\subset\ker(\pi)$. Since $g\pi$ has rank one, we have $\rho_{2}(g\pi)=0$ and $\rho_{1}(g\pi)=\frac{\|g\pi g\pi\|}{\|g\pi\|}>0$. Therefore, we have $\rho_{1}(g\pi)>\rho_{2}(g\pi)$ and $g\pi\in\overline{\Gamma}_{\nu}$, which contradicts (3). ∎
@@ 3.3. Construction of the Schottky measure
[…]
@@
Let $\Gamma$ be a measurable space, let $\mathbb{A}$ be a measurable binary relation on $\Gamma$ and let $10\leq\rho<$ 1. Let $\nu_{s}$ be a measure on $\Gamma$. We say that $\nu_{s}$ is $\rho$-Schottky for $\mathbb{A}$ if:
@@
Given $\Gamma$ be a measurable space, given $\mathbb{A}\subset\mathbb{A}^{\prime}$ two measurable binary relations on $\Gamma$, given $10\leq\rho\leq\rho^{\prime}<$ 1 and given $\nu$ be a measure on $\Gamma$, if $\nu_{s}$ is $\rho$-Schottky for $\mathbb{A}$, then it is also $\rho^{\prime}$-Schottky for $\mathbb{A}^{\prime}$.
@@
Such a $u_{m+1}$ exists because $\{u\in E\,|\,[u]\in\partial_{u}\Gamma\}\cup\{0\}$ is $\Gamma$-invariant by definition and it is not $\{0\}$ by Lemma 3.11. Hence $\{u\in E\,|\,[u]\in\partial_{u}\Gamma\}$ cannot be included in $\bigcup_{\begin{subarray}{c}i_{1}<\dots<i_{k}\end{subarray}}\langle u_{i_{j}}\rangle_{1\leq j\leq k}$, which is a finite union of hyperplanes. Then $[u_{m+1}]\in\partial\Gamma$ and we can easily check that $([u_{1}],\dots,[u_{m+1}])$ is in general position.
@@
Let $h\in\mathrm{End}(E)\setminus\{0\}$. Let $I^{h}:=\{i\,|\,u_{i}\in\ker(h)\}$. Since the family $([u_{1}],\dots,[u_{N}])$ is in general position and $\langle u_{i}\rangle_{i\in I}\subset\ker(h)$, we have #$\#I^{h}\leq d-1$. Let $J^{h}:=\{j\,|\,w_{j}\in\ker({}^{t}h)\}$. By the same argument, we have #$\#J^{h}\leq d-1$. For all $h\in\mathrm{End}(E)\setminus\{0\}$, let:
@@
Now let $h\in\mathrm{End}(E)\setminus\{0\}$. The set of indices $\{k\,|\,\pi_{k}\mathbb{A}^{\varepsilon}h\}$ is $\{k\,|\,w_{k}\mathbb{A}^{\varepsilon}h\}$, which has cardinality at least $N-d+1\geq(1-\rho)N$. Indeed, since $\psi(h)\geq\varepsilon$, a sufficient condition to have $w_{j}\mathbb{A}^{\varepsilon}h$ is for $j$ not to be included in a set $J$ that realises the maximum in the definition of $\psi(h)$. By the same argument #$\#\{k\,|\,h\mathbb{A}^{\varepsilon}\pi_{k}\}\geq N-d+1$. ∎
@@
Let $E$ be a Euclidean space, let $\nu$ be a strongly irreducible and proximal probability measure on $\mathrm{End}(E)$, let $10<\rho<$ 1 and let $f=(0,1)\to(0,1)$. There exist an integer $N$, two constants $\alpha,\varepsilon\in(0,1)$, an integer $m$ and a family $(S_{k})_{1\leq k\leq N}$ of measurable subsets of $\mathrm{End}(E)$ such that $\nu^{*m}(S_{k})\geq\alpha$ for all $1\leq k\leq N$ and such that:
@@
#$\#\{(i,j)\,|\,S^{\prime}_{i}\mathbb{A}^{\varepsilon}S^{\prime}_{k}\mathbb{A}^{\varepsilon}S^{\prime}_{i}\mathbb{A}^{\varepsilon}S^{\prime}_{j}\mathbb{A}^{\varepsilon}S^{\prime}_{k}\mathbb{A}^{\varepsilon}S^{\prime}_{j}\}\geq(1-2\rho)N(1-3\rho)N>0.$
[…]
@@
Let $E$ be a Euclidean space, let $\Gamma\in\{\mathrm{End}(E),\mathrm{GL}(E)\}$ and let $\nu$ be a strongly irreducible and proximal probability measure on $\Gamma$, let $10<\rho<$ 1 and let $\delta:(0,1)\to(0,1)$. There exists an integer $m$, two constants $\alpha,\varepsilon\in(0,1)$ and a probability measure $\tilde{\nu}_{s}$ on $\Gamma^{m}$ such that:
@@
    We have $\sigma_{*}\Pi_{*}\tilde{\nu}_{s}\left[0,\delta(\varepsilon)\right]=1$ i.e., for all $(s_{1},\dots,s_{m})$ in the support of $\tilde{\nu}_{s}$, we have $\sigma(s_{1}\cdots s_{m})\geq\delta(\varepsilon)$.
@@
[…]

### v3

@@ 2.2. Local-to-global properties for the alignment
Let $f,g,h$ be non-zero matrices such that the product fghfgh is well defined and let $0<\varepsilon\leq 1$. Assume that $\sigma(g)\leq\varepsilon^{2}/4$ and that $f\mathbb{A}^{\varepsilon}g\mathbb{A}^{\frac{\varepsilon}{2}}h$. Then $f\mathbb{A}^{\frac{\varepsilon}{2}}gh$.
@@
Let $f$, $g$ and $h$ be non-zero matrices such that the product fghfgh is well defined and let $0<\varepsilon\leq 1$. Assume that $f\mathbb{A}^{\varepsilon}g\mathbb{A}^{\varepsilon}h$ and that $\sigma(g)\leq\varepsilon^{4}/4$. Then $\sigma(fgh)\leq 4\varepsilon^{-4}\sigma(f)\sigma(g)\sigma(h)\geq 4\varepsilon^{-4}\sigma(g)$.
@@
Let $2\leq k\leq n-1$. First by Lemma 2.9 applied to $(g_{k},\dots,g_{n})$, we have $g_{k}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k+1}\cdots g_{n}$. Again by Lemma 2.9 applied to $(^{t}g_{k-1},\dots,^{t}g_{0})$, we have $g_{0}\cdots g_{k-2}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k-1}$. Moreover, we have $g_{k-1}\mathbb{A}^{\varepsilon}g_{k}$ an $\sigma(g_{k-1}),\sigma(g_{k})\leq\varepsilon^{2}/12$. Therefore, by Lemma 2.14, we have $g_{0}\cdots g_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}g_{k}\cdots g_{n}$. ∎
@@ 2.3. Link between singular values and eigenvalues
To get (34), we apply Corollary 2.13. We get a line $l^{\infty}$ such that for any sequence $u_{n}\in U^{1}(g^{n})\setminus\{0\}$, we have $u_{n}]\to l^{\infty}$. Let $l^{\infty}$ be such a line and $(u_{n})_{n\geq 0}$ be such a sequence. Moreover, for all $u\in U^{1}(g)$, we have $\mathrm{d}([u],l^{\infty})\leq\frac{2\sigma(g)}{\varepsilon}$. Now we only need to show that $l^{\infty}=E^{+}(g)$. Let $e\in E^{+}(g)$. Then we have $\|ge\|=\rho_{1}(g)\|e\|$ so by (32), $e\in V^{\frac{\varepsilon}{2}}(g)$. Moreover, $e$ is an eigenvector so $e\in\mathbb{K}ge$ and as a consequence $e\in U^{\frac{\varepsilon}{2}}(g)$. This reasoning holds for all powers of $g$ so we have $e\in U^{\frac{\varepsilon}{2}}(g^{n})$ for all $n$. Moreover, $\sigma(g^{n})\to 0$ by(33) so by Lemma 2.5, the projective diameter of $U^{\frac{\varepsilon}{2}}(g^{n})$ goes to zero so $u_{n}]\to[e]$ and therefore $l^{\infty}=[e]$. ∎
@@ 3. Random products and extractions
We use the following notations. Given $\Gamma$ a metric semi-group with a unit $\mathbf{1}_{\Gamma}$, we write $\widetilde{\Gamma}:=\bigsqcup_{n\geq 0}\Gamma^{n}$ for the set of words on the alphabet $\Gamma$ and write $\odot$ for the concatenation product. We write $()$ for the empty word.
@@
Given $(w_{k})_{k\geq 0}\in\mathbb{N}^{\mathbb{N}}$, we write $\overline{w}_{k}:=w_{0}+\cdots+w_{k-1}$ for all $k\in\mathbb{N}$ and given $(\gamma_{n})_{n\geq 0}\in\Gamma^{\mathbb{N}}$, we write $\gamma^{w}_{k}:=\gamma_{\overline{w}_{k}}\cdots\gamma_{\overline{w}_{k+1}-1}$ and $\widetilde{\gamma}^{w}_{k}:=\left(\gamma_{\overline{w}_{k}},\dots,\gamma_{\overline{w}_{k+1}-1}\right)$ for all $k\in\mathbb{N}$. That way, for all sequence $(w_{k})_{k\geq 0}\in\mathbb{N}^{N}N$ that is not stationary to 0 and for all sequence $(\gamma_{n})_{n\geq 0}\in\Gamma^{\mathbb{N}}$, we have $\bigodot_{k=0}^{+\infty}\widetilde{\gamma}^{w}_{k}=(\gamma_{n})_{n\geq 0}$.
@@ 3.2. Rank one boundary of a semi-group
$\partial_{u}\Gamma$ $:=\left\{[hv]\,\middle|\,[h]\in\partial\Gamma,\;v\in E\setminus\ker(h)\right\}\subset\mathrm{P}(E)$
$\partial_{w}\Gamma$ $:=\left\{[fh]\,\middle|\,[h]\in\partial\Gamma,\;f\in E^{*}\setminus\ker({{}^{t}h})\right\}\subset\mathrm{P}(E^{*}).$
@@
Note that the space $\mathrm{P}(\mathrm{End}(E))$ is metrizable so the closure is characterized by sequences. Let $\pi$ be a rank one endomorphism. Saying that $\pi]\in\partial\Gamma$ is equivalent to saying that there is a sequence $(\gamma_{n})\in\left(\Gamma\setminus\{0\}\right)^{\mathbb{N}}$ such that $\gamma_{n}]\to[\pi]$. Note also that the product map is continuous so $\overline{\Gamma}$ is a semi-group. Therefore, for all $\pi_{1}],[\pi_{2}]\in\partial\Gamma$, and for all $g\in\Gamma$ such that $\pi_{1}g\pi_{2}\neq 0$, we have $\pi_{1}g\pi_{2}]\in\partial\Gamma$.
Let $v_{1}\in\partial_{u}\Gamma$ and let $f_{2}\in\partial_{v}\Gamma$. Let $v_{2}\in E$ and $f_{1}\in E^{*}$ be such that $v_{1}f_{1}]\in\partial\Gamma$ and $v_{2}f_{2}]\in\partial\Gamma$. By definition of the irreducibility, there is an element $g\in\Gamma$ such that $f_{1}gv_{2}\neq 0$. Let $g$ be such an element. Then we have $v_{1}f_{1}gv_{2}f_{2}\neq 0$ hence $v_{1}f_{1}gv_{2}f_{2}]\in\partial\Gamma$. Moreover $v_{1}f_{1}gv_{2}f_{2}]=[v_{1}f_{2}]$ because $f_{1}gv_{2}$ is a scalar, therefore $v_{1}f_{2}]\in\partial\Gamma$. ∎
@@
Reciprocally, let $\pi]\in\partial\nu$. By irreducibility, there exists $g\in\overline{\Gamma}_{\nu}$ such that $g(\mathrm{im}(\pi))\not\subset\ker(\pi)$. Since $g\pi$ has rank one, we have $\rho_{2}(g\pi)=0$ and $\rho_{1}(g\pi)=\frac{\|g\pi g\pi\|}{\|g\pi\|}>0$. Therefore, we have $\rho_{1}(g\pi)>\rho_{2}(g\pi)$ and $g\pi\in\overline{\Gamma}_{\nu}$, which contradicts (3). ∎
@@ 3.3. Construction of the Schottky measure
In this paragraph $E$ is a vector space of dimension $d\geq 2$ over a local field $(\mathbb{K},|\cdot|)$. We endow $E$ with a norm $\|\cdot\|$ that is Euclidean when $\mathbb{K}=\mathbb{R}$, Hermitian when $\mathbb{K}=\mathbb{C}$ and ultra-metric otherwise. We endow $\mathrm{End}(E)$ with the operator norm. We recall the notation introduced in Definition 1.3. Given $0<\varepsilon\leq 1$ and $g,h$ two matrices, we write $g\mathbb{A}^{\varepsilon}h$ when $\|gh\|\geq\varepsilon\|g\|\|h\|$.
@@
Let $\Gamma$ be a measurable space, let $\mathbb{A}$ be a measurable binary relation on $\Gamma$ and let $0\leq\rho<1$. Let $\nu_{s}$ be a measure on $\Gamma$. We say that $\nu_{s}$ is $\rho$-Schottky for $\mathbb{A}$ if:
@@
Given $\Gamma$ be a measurable space, given $\mathbb{A}\subset\mathbb{A}^{\prime}$ two measurable binary relations on $\Gamma$, given $0\leq\rho\leq\rho^{\prime}<1$ and given $\nu$ be a measure on $\Gamma$, if $\nu_{s}$ is $\rho$-Schottky for $\mathbb{A}$, then it is also $\rho^{\prime}$-Schottky for $\mathbb{A}^{\prime}$.
@@
Such a $u_{m+1}$ exists because $\{u\in E\,|\,[u]\in\partial_{u}\Gamma\}\cup\{0\}$ is $\Gamma$-invariant by definition and it is not $\{0\}$ by Lemma 3.11. Hence $\{u\in E\,|\,[u]\in\partial_{u}\Gamma\}$ cannot be included in $\bigcup_{\begin{subarray}{c}i_{1}<\dots<i_{k}\end{subarray}}\langle u_{i_{j}}\rangle_{1\leq j\leq k}$, which is a finite union of hyperplanes. Then $u_{m+1}]\in\partial\Gamma$ and we can easily check that $([u_{1}],\dots,[u_{m+1}])$ is in general position.
@@
Let $h\in\mathrm{End}(E)\setminus\{0\}$. Let $I^{h}:=\{i\,|\,u_{i}\in\ker(h)\}$. Since the family $([u_{1}],\dots,[u_{N}])$ is in general position and $\langle u_{i}\rangle_{i\in I}\subset\ker(h)$, we have $I^{h}\leq d-1$. Let $J^{h}:=\{j\,|\,w_{j}\in\ker({}^{t}h)\}$. By the same argument, we have $J^{h}\leq d-1$. For all $h\in\mathrm{End}(E)\setminus\{0\}$, let:
@@
Now let $h\in\mathrm{End}(E)\setminus\{0\}$. The set of indices $\{k\,|\,\pi_{k}\mathbb{A}^{\varepsilon}h\}$ is $\{k\,|\,w_{k}\mathbb{A}^{\varepsilon}h\}$, which has cardinality at least $N-d+1\geq(1-\rho)N$. Indeed, since $\psi(h)\geq\varepsilon$, a sufficient condition to have $w_{j}\mathbb{A}^{\varepsilon}h$ is for $j$ not to be included in a set $J$ that realises the maximum in the definition of $\psi(h)$. By the same argument $\{k\,|\,h\mathbb{A}^{\varepsilon}\pi_{k}\}\geq N-d+1$. ∎
@@
Let $E$ be a Euclidean space, let $\nu$ be a strongly irreducible and proximal probability measure on $\mathrm{End}(E)$, let $0<\rho<1$ and let $f=(0,1)\to(0,1)$. There exist an integer $N$, two constants $\alpha,\varepsilon\in(0,1)$, an integer $m$ and a family $(S_{k})_{1\leq k\leq N}$ of measurable subsets of $\mathrm{End}(E)$ such that $\nu^{*m}(S_{k})\geq\alpha$ for all $1\leq k\leq N$ and such that:
@@
$\{(i,j)\,|\,S^{\prime}_{i}\mathbb{A}^{\varepsilon}S^{\prime}_{k}\mathbb{A}^{\varepsilon}S^{\prime}_{i}\mathbb{A}^{\varepsilon}S^{\prime}_{j}\mathbb{A}^{\varepsilon}S^{\prime}_{k}\mathbb{A}^{\varepsilon}S^{\prime}_{j}\}\geq(1-2\rho)N(1-3\rho)N>0.$
Hence the set of all possible values for $i_{k},j_{k}$ is non-empty. Let $m^{\prime}$ be the smallest common multiple of $\{n_{k}+n_{i_{k}}\,|\,1\leq k\leq N\}$, let $m^{\prime\prime}$ be the smallest common multiple of $\{n_{k}+n_{j_{k}}\,|\,1\leq k\leq N\}$ and let $m=m^{\prime}+m^{\prime\prime}$. Let $k\in\{1,\dots,N\}$. We define $p_{k}:=\frac{m^{\prime}}{n_{k}+n_{i_{k}}}$ and $q_{k}:=\frac{m^{\prime\prime}}{n_{k}+n_{j_{k}}}$ and:
@@
Let $E$ be a Euclidean, Hermitian or ultra-metric space, let $\Gamma\in\{\mathrm{End}(E),\mathrm{GL}(E)\}$ and let $\nu$ be a strongly irreducible and proximal probability measure on $\Gamma$, let $0<\rho<1$ and let $\delta:(0,1)\to(0,1)$. There exists an integer $m$, two constants $\alpha,\varepsilon\in(0,1)$ and a probability measure $\tilde{\nu}_{s}$ on $\Gamma^{m}$ such that:
@@
    We have $\sigma_{*}\Pi_{*}\tilde{\nu}_{s}\left[0,\delta(\varepsilon)\right]=1$ i.e., for all $(s_{1},\dots,s_{m})$ in the support of $\tilde{\nu}_{s}$, we have $\sigma(s_{1}\cdots s_{m})\leq\delta(\varepsilon)$.
@@
Then $f\leq\frac{N}{\alpha}$ because we assumed that $\nu^{*m}(S_{k})\geq\alpha$ for all $k$. Moreover, we can check that $\int_{\Gamma^{m}}fd\nu^{\otimes m}=N$ and for all $k\in\{1,\dots,N\}$ we have $\int_{\Pi^{-1}(S_{k})}fd\nu^{\otimes m}\geq 1$. Let:
$\tilde{\nu}_{s}:=\frac{f\nu^{\otimes m}}{\int_{\Gamma^{m}}fd\nu^{\otimes m}}\quad\text{and}\quad\nu_{s}:=\Pi_{*}\tilde{\nu}_{s}.$